# Stability for the 2-D plane Poiseuille flow in finite channel

### 31 Dec 2023 v1

@@ Abstract.
In this paper, we study the stability for 2-D plane Poiseuille flow $(1-y^{2},0)$ in a channel $\mathbb{T}\times(-1,1)$ with non-slip boundary condition. Three effects are studied in this paper: enhanced dissipation, inviscid damping and boundary layer effect. For the Navier-Stokes equations around Poiseuille flow with non-slip boundary condition, we prove that if the initial perturbation of velocity $u_{0}$ satisfies
$\|u_{0}\|_{H^{2+}}\leq\epsilon_{0}\nu^{\frac{3}{4}}$
for some constants $\epsilon_{0}>0$ independent of the viscosity $\nu$, then the solution to 2-D Navier-Stokes equations does not transit from plane Poiseuille flow. Meanwhile, for the Navier-slip boundary problem, we prove that if the initial perturbation of velocity $u_{0}$ satisfies
$\|u_{0}\|_{H^{\frac{3}{2}+}}\leq\epsilon_{1}\nu^{\frac{1}{2}}$
for some constants $\epsilon_{1}>0$ independent of the viscosity $\nu$, then the solution to 2-D Navier-Stokes equations does not transit from plane Poiseuille flow, which improves the result of [17] from $\nu^{\frac{3}{4}}$ to $\nu^{\frac{1}{2}}$.
@@ 1. Introduction
In this paper, we consider the 2-D incompressible Navier-Stokes equations in a channel $\Omega:=\mathbb{T}\times(-1,1)$:
@@
It is necessary to impose the boundary condition for (1.2). In this paper, we consider two types of boundary conditions:
Case I. non-slip boundary condition:
$u(t;x,\pm 1)=0;$ (1.3)
and
Case II. Navier-slip boundary condition:
@@
where $\Phi$ is the stream function and $\omega=\partial_{y}u_{1}-\partial_{x}u_{2}$ is the vorticity. In addition, the boundary conditions (1.3) and (1.4) are rewritten as follows respectively
$(\Phi,\partial_{y}\Phi)(t;x,\pm 1)=0$ (1.6)
and
@@
Our interest in this paper is to understand the long time dynamics behavior of (1.5) with (1.6), and (1.5) with (1.7) respectively for $0<\nu\ll 1$.
@@
To understand the nonlinear dynamics stability of shear flow, it is necessary to study the transition threshold problem mentioned before. There are lots of works devoted to determining the exponent $\gamma$ for different shear flows via the numerical or asymptotic analysis, see [15, 31, 36] and references therein. Recently, there are many progresses about the rigorous mathematical analysis on the transition threshold problem for different shear flows, including the Couette flow and Kolmogorov flow. For 3-D plane Couette flow, either $X$ is Gevrey class or Sobolev space, the transition threshold satisfies $\gamma\leq 1$ for the case without boundaries, see [2, 3, 4, 47]. The result $\gamma\leq 1$ also holds in a finite channel with non-slip boundary condition [12]. The transition threshold for 2-D Couette flow is proved as $\gamma\leq\frac{1}{2}$ in a channel with non-slip boundary condition [10], and $\gamma\leq\frac{1}{3}$ in $\mathbb{T}\times\mathbb{R}$ [32, 44]. More interesting related results can be found in [1, 5, 7, 20, 42].
@@
The first main result is stated as follows.
@@
Suppose that $\mathrm{div}\ u_{0}=0$ and $\|u_{0}\|_{H^{2+}}\leq\epsilon_{0}\nu^{3/4}$ for some constants $\epsilon_{0}$ independent of $\nu$, then the solution to (1.5) with (1.6) is global in time with
$\sum_{k\in\mathbb{Z}}\mathcal{F}_{k}\leq C\nu^{3/4},$ (1.8)
where the energy functional $\mathcal{F}_{k}$ is defined as
$\mathcal{F}_{k}:=\left\{\begin{array}[]{lll}\|(1-y^{2})^{\frac{1}{2}}\omega_{k}\|_{L^{\infty}L^{2}}+(\nu|k|)^{\frac{1}{4}}\|\omega_{k}\|_{L^{2}L^{2}}\\ \qquad\qquad\qquad+|k|^{\frac{1}{2}}(\|u_{k}\|_{L^{2}L^{2}}+\|u_{k}\|_{L^{\infty}L^{2}}),&k\not=0,\\ \|\overline{\omega}\|_{L^{\infty}L^{2}},&k=0.\end{array}\right.$ (1.9)
Remark 1.1.
Let us recall that in [17], the transition threshold with $\gamma\leq\frac{3}{4}$ has been proved for the Navier-slip boundary value problem. Our Theorem 1.1 obtain the same transition threshold with $\gamma\leq\frac{3}{4}$ for the case of non-slip boundary value problem. This result implies that the boundary layer effect induced by non-slip boundary condition is not the main effect resulting in the strong instability. It is interesting to prove the stability threshold $\gamma\leq\frac{1}{2}$ mentioned in [9].
Remark 1.2.
To our knowledge, the Theorem 1.1 is the first result trying to answer the problem proposed in [9], which focuses on the transition threshold for 2-D plane Poiseuille flow with non-slip boundary condition. In our work, the stability mechanisms (including enhanced dissipation, inviscid damping) play the key role. The result $\gamma\leq\frac{3}{4}$ can be viewed as the consequence of the balance between the two good effect and boundary layer effect.
Remark 1.3.
In the energy (1.9), the weight $(1-y^{2})^{\frac{1}{2}}$ in $\|(1-y^{2})^{\frac{1}{2}}\omega_{k}\|_{L^{\infty}L^{2}}$ is introduced to correct the singularity near boundaries $\{y=\pm 1\}$, which is due to the boundary layer effect resulted from the non-slip boundary condition. And the $|k|^{\frac{1}{2}}\|u_{k}\|_{L^{2}L^{2}}$ is due to the inviscid damping effect of Poiseuille flow.
The second result is to improve the transition threshold of Navier-slip boundary value problem studied in [17]. Precisely, we improve the result of [17] from $\nu^{3/4}$ to $\nu^{1/2}$, which is stated as the following theorem.
Theorem 1.2.
Suppose that $\mathrm{div}u_{0}=0,u_{0}\in H^{\frac{3}{2}+}$ with $\|u_{0}\|_{H^{\frac{3}{2}+}}\leq\epsilon_{1}\nu^{1/2}$ for some suitable small $0<\epsilon_{1}\ll 1$ independent of $\nu$, then (1.5) with (1.4) admits a global in time solution $\omega$ with
$\sum_{k\in\mathbb{Z}}\mathcal{E}_{k}\leq C\nu^{1/2},$ (1.10)
@@
$\mathcal{E}_{k}:=\left\{\begin{array}[]{lll}\|\omega_{k}\|_{L^{\infty}L^{2}}+(\nu|k|^{2})^{1/6}\|\omega_{k}\|_{L^{2}L^{2}}+|k|^{1/2}\|u_{k}\|_{L^{2}L^{2}},&k\not=0,\\ \|\overline{\omega}\|_{L^{\infty}L^{2}},&k=0.\end{array}\right.$ (1.11)
@@
$\sum_{k\in\mathbb{Z}}\|u_{k}\|_{L^{\infty}L^{\infty}}\leq C\nu^{\frac{1}{2}},$
@@
Even for the Navier-slip boundary value problem, it is not sure that the transition threshold $\gamma\leq\frac{1}{2}$ is optimal or not. It is very interesting and challenging to verify the sharpness of $\frac{1}{2}$.
Let us give some comments about our results.
To our knowledge, our Theorem 1.1 firstly gives the transition threshold for the 2-D plane Poiseuille flow with non-slip boundary condition, i.e., $\gamma\leq\frac{3}{4}$. Compared with [17], our result implies that the boundary layer effect induced by non-slip boundary condition is not the main mechanism resulting in the strong nonlinear instability. The linear effects are formulated in the functional (1.9), including the enhanced dissipation, inviscid damping and boundary layer effect. The key point is the space-time estimates for linearized Navier-Stokes system with source terms, which corresponding to the careful resolvent estimates of Orr-Sommerfeld equation around plane Poiseuille flow established in [9].
In addition, compared with [17], the new trouble in this paper is boundary layer effect. Note that near boundaries, it holds that $1-y^{2}\sim 1-|y|$. To control the boundary layer effect, the weight estimate with weight $1-y^{2}$ is needed, which avoid the loss of $\nu$ in $L^{\infty}L^{2}$-estimate for vorticity. If one only considers the $L^{\infty}L^{2}$-estimate for vorticity without weight $1-y^{2}$, there is loss of $\nu^{\frac{1}{4}}$ in the estimate, which is resulted from the boundary layer effect. In addition, to achieve the transition threshold $\gamma\leq\frac{3}{4}$ as in [17], the enhanced dissipation and inviscid damping effect are needed, which contribute the stabilizing mechanism to balance the boundary layer effect.
Our second result improves the result of [17] from $\nu^{3/4}$ to $\nu^{1/2}$. Compared with [17], the main contribution to the improvement is the inviscid damping effect. Indeed, the result $\nu^{3/4}$ in [17] is due to the enhanced dissipation effect. To improve it to $\nu^{1/2}$, the key point is the inviscid damping effect and the good structure of the Rayleigh part (or inviscid part)
$(1-y^{2})+2\Delta^{-1},$
or in the Fourier version, namely,
$(1-y^{2})+2(\partial_{y}^{2}-k^{2})^{-1}.$
The above Rayleigh structure gives good positive estimates for velocity and vorticity(with weight), which along with the diffusion yield better space-time $L^{2}L^{2}$ estimate to vorticity, see Section 2 for instance. This implies the good stability mechanism of Poiseuille flow in some sense.
Moreover, with the good positive structure, one may notice that our proof is different from that of [17] and [14, 18]. The proof for Theorem 1.2 relies on a direct energy argument with some positive structure for the toy model around Poiseuille flow, see Section 2 for formal analysis.
@@ 2. Sketch and key ideas of the proof
2.1. The stability for Poiseuille flow with non-slip boundary condition
To obtain the transition threshold of plane Poiseuille flow with non-slip boundary condition, the key point is to establish the space-time estimates for the following linearized system
$\left\{\begin{array}[]{lll}\partial_{t}\omega-\nu(\partial_{y}^{2}-k^{2})\omega+ik(1-y^{2})\omega+2ik\varphi=-ikf_{1}-\partial_{y}f_{2},\\ (\partial_{y}^{2}-k^{2})\varphi=\omega,\ \ u=(\partial_{y},-ik)\varphi,\\ \varphi(t,k,\pm 1)=\partial_{y}\varphi(t,k,\pm 1)=0,\\ \omega|_{t=0}=\omega_{0}(k,y),\end{array}\right.$
where $f_{i}\in L^{2}$ and $k\not=0$. The basic space-time estimates are the $L^{2}L^{2}$ estimates for vorticity and velocity field, which are due to enhanced dissipation and inviscid damping induced by plane Poiseuille flow. See Section 3 and [9] for details.
Compared with the Navier-slip boundary condition in [17], the new effect may result in instability is the boundary layer effect, which is resulted from the non-slip boundary condition. Precisely, due to the singularity occurring near the boundaries, there is loss of $\nu$ in the space-time estimates if following the arguments of [17], which will lead to a threshold worse than $\nu^{3/4}$. To overcome the difficulty, the main idea is to introduce the suitable weight to control this effect near boundaries $\{y=\pm 1\}$. Since near $\{y=\pm 1\}$, it holds that $1-y^{2}\sim 1-|y|$, the suitable weight in our argument is chosen as $1-y^{2}$. Precisely, we deduce the weighed estimate as follows:
$\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{\infty}L^{2}}^{2}+\nu\|(1-y^{2})^{\frac{1}{2}}(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}+|k|\|(1-y^{2})\omega\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $|k|^{-1}\|(\partial_{y},|k|)\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2},$
see Section 3 for details. There are some advantages about this weight: it corrects the behavior of the vorticity near boundaries and it can be viewed as the good weight in some Hardy-type inequalities. If we consider the estimates without the weight, we only obtain that
$\nu^{\frac{1}{2}}\|\omega\|_{L^{\infty}L^{2}}^{2}+\nu^{\frac{3}{2}}\|(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $|k|^{-1}\|(\partial_{y},|k|)\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|(f_{1},f_{2})\|_{L^{2}L^{2}}^{2},$
There is loss of $\nu$ in the above estimates, and the weight $1-y^{2}$ avoid this loss of $\nu$.
With the above arguments, the space-time estimates for nonlinear stability read as
$\mathcal{F}_{k}\lesssim\|\omega_{0}\|_{H^{2+}}+\mathrm{nonlinear\ terms},$
where the energy functional for $k\not=0$ is defined as
$\mathcal{F}_{k}:=\|(1-y^{2})^{\frac{1}{2}}\omega_{k}\|_{L^{\infty}L^{2}}+(\nu|k|)^{\frac{1}{4}}\|\omega_{k}\|_{L^{2}L^{2}}+|k|^{\frac{1}{2}}(\|u_{k}\|_{L^{2}L^{2}}+\|u_{k}\|_{L^{\infty}L^{2}}).$
The term $\|(1-y^{2})^{\frac{1}{2}}\omega_{k}\|_{L^{\infty}L^{2}}$ is due to boundary layer effect, $(\nu|k|)^{\frac{1}{4}}\|\omega_{k}\|_{L^{2}L^{2}}$ is due to enhanced dissipation, and $|k|^{\frac{1}{2}}\|u_{k}\|_{L^{2}L^{2}}$ is due to inviscid damping. In other words, the result is the consequence of the balance between the enhanced dissipation, inviscid damping and boundary layer effect.
Then the nonlinear stability follows from continuous argument, see Subsection 5.1 for details.
2.2. The improved transition threshold with Navier-slip boundary condition
@@
$\left\{\begin{array}[]{lll}\partial_{t}\omega-\nu(\partial_{y}^{2}-k^{2})\omega+ik(1-y^{2})\omega+2ik\varphi=-ikf_{1}-\partial_{y}f_{2},\\ (\partial_{y}^{2}-k^{2})\varphi=\omega,\ u=(\partial_{y},-ik)\varphi,\\ (\varphi,\omega)|_{y=\pm 1}=(0,0),\ \omega|_{t=0}=\omega_{0}(k,y),\end{array}\right.$ (2.1)
where $\omega_{0}(k,\pm 1)=0$ and $k\not=0$.
@@
To establish the improved space-time estimates, we consider the energy form as follows
$\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}$ $\|\omega\|_{L^{2}}^{2}+\nu\|(\partial_{y},|k|)\omega\|_{L^{2}}^{2}$ (2.2)
$+ik\langle(1-y^{2})\omega+2\varphi,\omega\rangle=\langle-ikf_{1}-\partial_{y}f_{2},\omega\rangle:=\mathrm{RHS}.$
The first key observation is the inviscid Rayleigh part (imaginary part) of (2.2) will give the space-time estimate for velocity field, which corresponds to the inviscid damping effect (integral form) of velocity field. More precisely, the Rayleigh part yields the following positive estimate (see Lemma 4.2)
$\|(\varphi^{\prime},|k|\varphi)\|_{L^{2}}^{2}+|k|^{2}\|\varphi\|_{L^{2}}^{2}\lesssim\langle(1-y^{2})w+2\varphi,w\rangle,$
which implies the following improved estimate
$\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}+\left(\|\varphi^{\prime}\|_{L^{2}}^{2}+|k|^{2}\|\varphi\|_{L^{2}}^{2}\right)$ (2.3)
$\lesssim$ $\mathrm{Re}\langle(1-y^{2})\omega+2\varphi,\omega\rangle\lesssim|k|^{-1}|\mathrm{Im}\ \mathrm{RHS}|,$
which implies the estimate for velocity field. See the proof of Proposition 4.2 for details.
It should be pointed out that the estimate (2.3) is resulted from the Rayleigh part and it is valid uniform in $\nu$, and this implies the intrinsic structure and stabilization mechanism of the Poiseuille flow.
The second key observation is the improvement of the $L^{2}L^{2}$-estimate for $\omega$. Precisely, with (2.3) at hands, by applying Lemma 4.1, we have
$(\nu|k|^{2})^{1/3}\|\omega\|_{L^{2}}^{2}\lesssim(\nu|k|^{2})^{1/3}\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{4/3}\|\partial_{y}\omega\|_{L^{2}}^{2/3}\lesssim\nu\|\partial_{y}\omega\|_{L^{2}}^{2}+|k|\|(1-y^{2})^{\frac{1}{2}}\omega\|_{L^{2}}^{2}$
for $\omega|_{y=\pm 1}=0$. Then we establish
$\frac{\mathrm{d}}{\mathrm{d}t}$ $\|\omega\|_{L^{2}}^{2}+c(\nu|k|^{2})^{1/3}\|\omega\|_{L^{2}}^{2}+\cdots\lesssim|\mathrm{RHS}|,$ (2.4)
which suggests that the $L^{2}L^{2}$-estimate for $\omega$ in [17] can be improved as
$\|e^{c_{0}(\nu|k|^{2})^{1/3}t}\omega\|_{L^{\infty}L^{2}}^{2}+(\nu|k|^{2})^{1/3}\|e^{c_{0}(\nu|k|^{2})^{1/3}t}$ $\omega\|_{L^{2}L^{2}}^{2}+\cdots$ (2.5)
$\lesssim\|\omega_{0}\|_{L^{2}}^{2}+\mathrm{source\ terms}.$
Therefore, the improved space-time estimates for toy model (2.1) follow. See Proposition 4.2 for details.
The nonlinear stability and transition threshold are also based on the improved space-time estimates in Proposition 4.2. Indeed, in [17], we only used the enhanced dissipation effect to obtain that transition threshold $\nu^{\gamma}$ with $\gamma\leq 3/4$. In this paper, additional to the enhanced dissipation, the inviscid damping for velocity field is established. With the effects and Proposition 4.2, the space-time estimates of the $x$-nonzero modes (i.e., in the Fourier version with $k\not=0$) for nonlinear problem read as
$\mathcal{E}_{k}\lesssim\|\omega_{0,k}\|_{L^{2}}+\mathrm{nonlinear\ terms},$
where the energy functional $\mathcal{E}_{k}$ with $k\not=0$ is defined as
$\mathcal{E}_{k}=\|\omega_{k}\|_{L^{\infty}L^{2}}+(\nu|k|^{2})^{1/6}\|\omega_{k}\|_{L^{2}L^{2}}+|k|^{1/2}\|u_{k}\|_{L^{2}L^{2}}.$
Compared with [17], the term $(\nu|k|^{2})^{1/6}\|\omega_{k}\|_{L^{2}L^{2}}$ is due to the improved space-time estimates, see the formal analysis (2.4)-(2.5). If following the arguments of [17], we can only obtain the bound of $(\nu|k|^{2})^{1/4}\|\omega_{k}\|_{L^{2}L^{2}}$, which is worse than the bound obtained in this paper. In addition, the term $|k|^{1/2}\|u_{k}\|_{L^{2}L^{2}}$ is due to the inviscid damping. It should be noted that it is independent of $\nu$ and it would avoid the loss of $\nu$ in nonlinear interaction. See Subsection 5.2 for details.
With the above key points, one can improve the result of [17] from $\nu^{3/4}$ to $\nu^{1/2}$.

### 03 Mar 2024 v2

@@ Abstract.
In this paper, we study the stability for 2-D plane Poiseuille flow $(1-y^{2},0)$ in a channel $\mathbb{T}\times(-1,1)$ with Navier-slip boundary condition. We prove that if the initial perturbation for velocity field $u_{0}$ satisfies that $\|u_{0}\|_{H^{\frac{7}{2}+}}\leq\epsilon_{1}\nu^{2/3}$ for some suitable small $0<\epsilon_{1}\ll 1$ independent of viscosity coefficient $\nu$, then the solution to the Navier-Stokes equations is global in time and does not transit from the plane Poiseuille flow. This result improves the result of [18] from $3/4$ to $2/3$.
@@ 1. Introduction
In this paper, we consider the 2-D incompressible Navier-Stokes (N-S) equations in a channel $\Omega:=\mathbb{T}\times(-1,1)$:
@@
It is necessary to impose the boundary condition for (1.2). In this paper, we consider the Navier-slip boundary condition
@@
where $\Phi$ is the stream function and $\omega=\partial_{y}u_{1}-\partial_{x}u_{2}$ is the vorticity. In addition, the boundary condition (1.3) is rewritten as follows
@@
Our interest in this paper is to understand the long time dynamics behavior of (1.4) with (1.5) for $0<\nu\ll 1$.
@@
To understand the nonlinear dynamics stability of shear flow, it is necessary to study the transition threshold problem mentioned before. There are lots of works devoted to determining the exponent $\gamma$ for different shear flows via the numerical or asymptotic analysis, see [16, 32, 37] and references therein. Recently, there are many progresses about the rigorous mathematical analysis on the transition threshold problem for different shear flows, including the Couette flow and Kolmogorov flow. For 3-D plane Couette flow, either $X$ is Gevrey class or Sobolev space, the transition threshold satisfies $\gamma\leq 1$ for the case without boundaries, see [2, 3, 4, 48]. The result $\gamma\leq 1$ also holds in a finite channel with non-slip boundary condition [12]. The transition threshold for 2-D Couette flow is proved as $\gamma\leq\frac{1}{2}$ in a channel with non-slip boundary condition [10], and $\gamma\leq\frac{1}{3}$ in $\mathbb{T}\times\mathbb{R}$ [33, 45]. Recently, Coti Zelati et al. showed that the transition threshold for Couette flow in 3-D Bounssinesq equations enjoys $\gamma<1$ [15]. More interesting related results can be found in [1, 5, 7, 21, 43].
@@
The main result is stated as follows, which improves the transition threshold of Navier-slip boundary value problem studied in [18]. Precisely, we improve the result of [18] from $\nu^{3/4}$ to $\nu^{2/3}$, which is stated as the following theorem.
@@
Suppose that $\mathrm{div}u_{0}=0,u_{0}\in H^{\frac{7}{2}+}$ with $\|u_{0}\|_{H^{\frac{7}{2}+}}\leq\epsilon_{1}\nu^{2/3}$ for some suitable small $0<\epsilon_{1}\ll 1$ independent of $\nu$, then (1.4) with (1.3) admits a global in time solution $\omega$ with
$\sum_{k\in\mathbb{Z}}\mathcal{E}_{k}\leq C\nu^{2/3},$ (1.6)
@@
$\mathcal{E}_{k}:=\left\{\begin{array}[]{lll}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{\infty}L^{2}}+\nu^{\frac{1}{2}}|k|\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{2}L^{2}}\\ \qquad+(\nu|k|)^{1/4}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{2}L^{2}}+|k|^{1/2}\|e^{c\nu^{\frac{1}{2}}t}u_{k}\|_{L^{2}L^{2}},&k\not=0,\\ \|\overline{\omega}\|_{L^{\infty}L^{2}},&k=0.\end{array}\right.$ (1.7)
@@
$\sum_{k\in\mathbb{Z}}\|e^{c\nu^{\frac{1}{2}}t}u_{k}\|_{L^{\infty}L^{\infty}}\leq C\nu^{2/3},$
@@
Even for the Navier-slip boundary value problem, the transition threshold $\gamma\leq 2/3$ should not be optimal. It is very interesting and challenging to study the optimal transition threshold of the Poiseuille flow.
@@ 2. Sketch and key ideas of the proof
$\left\{\begin{array}[]{lll}\partial_{t}\omega+\mathcal{L}_{k}\omega=-ikf_{1}-\partial_{y}f_{2}-f_{3}-f_{4},\\ (\partial_{y}^{2}-k^{2})\varphi=\omega,\ u=(\partial_{y},-ik)\varphi,\\ (\varphi,\omega)|_{y=\pm 1}=(0,0),\ \omega|_{t=0}=\omega_{0}(k,y),\end{array}\right.$ (2.1)
where $\omega_{0}(k,\pm 1)=0$ and $k\not=0$, the linearized operator is defined as
$\mathcal{L}_{k}:=-\nu(\partial_{y}^{2}-k^{2})+ik(1-y^{2})+2ik(\partial_{y}^{2}-|k|^{2})^{-1}.$
@@
To improve the result of [18] to $\gamma\leq 2/3$, the inviscid damping effect is added in the space-time estimates in this paper. More precisely, the following space-time estimates are used in this paper
$\|e^{c\nu^{\frac{1}{2}}t}\omega\|_{L^{\infty}L^{2}}^{2}+\nu\|e^{c\nu^{\frac{1}{2}}t}(\partial_{y},|k|)\omega\|_{L^{2}L^{2}}^{2}$
$+(\nu|k|)^{\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}\omega\|_{L^{2}L^{2}}^{2}+|k|\|e^{c\nu^{\frac{1}{2}}t}u\|_{L^{2}L^{2}}^{2}$
$\lesssim$ $\|\Delta_{k}\omega_{0}\|_{L^{2}}^{2}+\nu^{-1}\|e^{c\nu^{\frac{1}{2}}t}(f_{1},f_{2})\|_{L^{2}L^{2}}^{2}+|k|^{-1}\|e^{c\nu^{\frac{1}{2}}t}(\partial_{y},|k|)f_{3}\|_{L^{2}L^{2}}^{2}$
$+\min\{(\nu|k|)^{-\frac{1}{2}},(\nu|k|^{2})^{-1}\}\|e^{c\nu^{\frac{1}{2}}t}f_{4}\|_{L^{2}L^{2}}^{2}.$
Compared with the space-time estimates in [18], in the above estimates, the part $|k|\|e^{c\nu^{\frac{1}{2}}t}u\|_{L^{2}L^{2}}^{2}$ is due to the inviscid damping effect of the Poiseuille flow.
To achieve the desired space-time estimates, following the idea of [9], the key point is to establish the resolvent estimates for the linearized operator around the Poiseuille flow, see Section 3 for details. With the resolvent estimates at hands, one can achieve the space-time estimates for the linearized N-S system.
For the nonlinear problem, the energy estimates for non-zero modes read as
$\mathcal{E}_{k}\lesssim\|\Delta_{k}\omega_{0,k}\|_{L^{2}}+\nu^{-\frac{3}{8}}|k|^{\frac{3}{8}}\|e^{c\nu^{\frac{1}{2}}t}(f_{1})_{k}\|_{L^{2}L^{2}}+\nu^{-\frac{1}{2}}\|e^{c\nu^{\frac{1}{2}}t}(f_{2})_{k}\|_{L^{2}L^{2}},$
where $\mathcal{E}_{k}$ is defined as
$\mathcal{E}_{k}:=$ $\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{\infty}L^{2}}+\nu^{\frac{1}{2}}|k|\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{2}L^{2}}$
$+(\nu|k|)^{1/4}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{2}L^{2}}+|k|^{1/2}\|e^{c\nu^{\frac{1}{2}}t}u_{k}\|_{L^{2}L^{2}}.$
In the above functional, the terms $\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{\infty}L^{2}}$ and $\nu^{\frac{1}{2}}|k|\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{2}L^{2}}$ are corresponding to the basic energy and heat diffusion, the term $(\nu|k|)^{1/4}\|e^{c\nu^{\frac{1}{2}}t}\omega_{k}\|_{L^{2}L^{2}}$ is due to the enhanced dissipation, and $|k|^{1/2}\|e^{c\nu^{\frac{1}{2}}t}u_{k}\|_{L^{2}L^{2}}$ is due to the inviscid damping effect of Poiseuille flow.
With the nonlinear estimates, the nonlinear stability follows from a continuous argument.
3. Resolvent estimates for Orr-Sommerfeld equation
In this section, we study the resolvent estimates for the Orr-Sommerfeld equation around the plane Poiseuille flow as follows
$\left\{\begin{array}[]{lll}-\nu(\partial_{y}^{2}-|k|^{2})w+ik[(1-y^{2}-\lambda)w+2\varphi]=F,\\ (\partial_{y}^{2}-|k|^{2})\varphi=w,\ \varphi(\pm 1)=w(\pm 1)=0,\end{array}\right.$ (3.1)
where $k\neq 0$ and $F\in H^{1}_{0}$.
The resolvent estimates for (3.1) are stated as follows.
Proposition 3.1.
Let $w$ be the solution to (3.1), then it holds that
$\nu^{\frac{3}{8}}$ $|k|^{\frac{5}{8}}(\|w\|_{L^{1}}+|k|^{\frac{1}{2}}\|u\|_{L^{2}})$ (3.2)
$+(\nu|k|)^{\frac{1}{2}}\|w\|_{L^{2}}+\nu^{\frac{3}{4}}|k|^{\frac{1}{4}}\|(\partial_{y},|k|)w\|_{L^{2}}\lesssim\|F\|_{L^{2}},$
$\nu^{\frac{3}{4}}$ $|k|^{\frac{1}{4}}\|w\|_{L^{2}}+\nu\|(\partial_{y},|k|)w\|_{L^{2}}+(\nu|k|)^{\frac{1}{2}}\|u\|_{L^{2}}\lesssim\|F\|_{H^{-1}_{k}},$
||| $\nu/k|^{\frac{1}{2}}\|(\partial_{y},|k|)w\|_{L^{2}}+|\nu/k|^{\frac{1}{4}}\|w\|_{L^{2}}$
$+|\nu/k|^{\frac{1}{8}}\|u\|_{L^{\infty}}+\|u\|_{L^{2}}\lesssim|k|^{-1}\|(\partial_{y},|k|)F\|_{L^{2}}.$
Moreover, there holds that
$\nu^{\frac{1}{6}}$ $|k|^{\frac{5}{6}}(|\lambda-1|^{\frac{1}{2}}+|\nu/k|^{\frac{1}{4}})^{\frac{1}{3}}\|u\|_{L^{2}}$ (3.3)
$+\nu^{\frac{2}{3}}|k|^{\frac{1}{3}}(|\lambda-1|^{\frac{1}{2}}+|\nu/k|^{\frac{1}{4}})^{\frac{1}{3}}\|(\partial_{y},|k|)w\|_{L^{2}}$
$+\nu^{\frac{1}{3}}|k|^{\frac{2}{3}}(|\lambda-1|^{\frac{1}{2}}+|\nu/k|^{\frac{1}{4}})^{\frac{2}{3}}\|w\|_{L^{2}}\lesssim\|F\|_{L^{2}},$
$\nu^{\frac{2}{3}}$ $|k|^{\frac{1}{3}}(|\lambda-1|^{\frac{1}{2}}+|\nu/k|^{\frac{1}{4}})^{\frac{1}{3}}\|w\|_{L^{2}}\lesssim\|F\|_{H^{-1}_{k}}.$
Here, $\|F\|_{H^{-1}_{k}}:=\inf\limits_{\{g\in H^{1}_{0}:\|g\|_{H^{1}_{k}}=1\}}\langle f,g\rangle$ with $\|g\|_{H^{1}_{k}}=\|(\partial_{y},|k|)g\|_{L^{2}}.$
Proof.
See [9]. ∎